# Transonic accretion flow in the mini discs of a binary black hole system

### v1

@@ Abstract
We study the general relativistic transonic accretion flow around the primary black hole, which forms the circumprimary disc (CPD), within a BBH system. The BBH spacetime is characterized by the mass ratio ($q$) and the separation distance ($z_{2}$) between the two black holes. We numerically solve the radial momentum and energy equations to obtain the accretion solutions. It is observed that the CPD can exhibit shock solutions, which exist for a wide range parameter space spanned by flow angular momentum ($\lambda$) and energy ($E$). We find that the shock parameter space is modified by $q$ and $z_{2}$. Investigations show that $q$ and $z_{2}$ also affect various shock properties, such as density compression and temperature compression across the shock fronts. Moreover, we calculate the spectral energy distributions (SEDs) of the CPD and examine how the SEDs are modified by $q$ and $z_{2}$ for both shock-free and shock-induced accretion solutions. SED is found to be nearly independent of the binary parameters. We depict that the mutual interaction between two black holes in a binary system hardly alters their individual accretion features. This study replicates the existing predictions in a different context of binary accretion.
accretion; accretion discs; black hole physics; hydrodynamics; shock waves
@@ I Introduction
Binary black hole (BBH) systems are powerful multi-messenger astrophysical sources of gravitational waves (GWs) and electromagnetic (EM) waves. The formation scenarios of BBHs are not clearly understood yet. In this regard, various astrophysical channels have been proposed in the literature. Notable studies by [1] and [2] indicate that supermassive black hole binaries (SMBHBs) form as a result of galaxy mergers. On the other hand, stellar-mass BBHs are generated through dynamic encounters in the environments of dense star clusters [3, 4]. Recently, it has been demonstrated that the gaseous dynamical friction can form a BBH system in the active galactic nucleus (AGN) discs [5].
@@
In addition to the understanding of the formation channels of BBHs, detecting them through observations is also a vital aspect in astrophysics. Meanwhile, LIGO and Virgo collaborations reveal that the gravitational interactions within BBHs can produce GWs [6], as predicted by general relativity, thereby providing direct evidence of stellar-mass BBHs through GWs. On the other hand, SMBHBs at sub-parsec separations are the strong GW sources [7], which may be detectable by the Pulsar Timing Arrays [8] or the future mission Laser Interferometer Space Antenna [9, 10]. Like AGNs, BBHs can radiate across the electromagnetic spectrum due to the accretion of matter from the surroundings. Therefore, in addition to GWs, analyzing EM signature from BBHs is also crucial for their study. In recent years, several EM observational works, including both the thermal and non-thermal emissions, have come out for the BBHs [11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Indeed, these investigations provide the valuable information about the binary evolution, mass accretion rate, disc inclination angle, mass and the spin of the individual black holes, etc.
@@
The accretion flow onto a BBH system is qualitatively different from a single black hole accretion. For BBH accretion, a truncated disc surrounds the entire binary system [21], called the circumbinary disc (CBD), where the matter flows inward due to the gravity of BBHs and the angular momentum transport through the viscous stress. Moreover, the tidal interaction between BBHs and CBD creates a cavity within the CBD, which defines its inner boundary. A significant fraction of the matter at this inner boundary falls toward the binary system and begins to orbit the individual black holes, resulting in the formation of a pair of mini discs. The mini discs surrounding the primary (more massive) and secondary (less massive) black holes are referred to as circumprimary discs (CPDs) and circumsecondary discs (CSDs), respectively. The presence of multiple discs (i.e., CBD, CPD, and CSD) complicates the study of BBH accretion compared to the accretion of individual black holes. Numerous theoretical models on BBHs accretion flow have been proposed in Newtonian gravity and general relativity (GR) [22, 23, 24, 25, 18, 19, 26, 21, and references therein]. All these models based on different physical conditions and explained the dynamics of CBD, mini discs, role of magnetic field, outflows or jets and EM spectral signature, etc., which indeed help to distinguish them from the ordinary AGNs. However, to the best of our knowledge, no body has conveyed the transonic acccretion flow for the BBH systems, where flow must satisfied the inner boundary condition set by the horizon [27, 28]. Toward this, for the first time we study the GR hydrodynamics of transonic accretion flow in the background of a BBH metric [29], which represents a system of two Schwarzschild black holes in an external gravitational field. We expect this study may provide the better understanding of the BBHs.
@@
Finding an exact analytical solution for binary or multiple black hole sources in GR is very challenging. Several attempts have been made to address this problem [30, 31], where it often requires the extra matter fields, commonly interpreted as struts or cosmological strings, to support the gravitational attraction between the black holes. However, all of them are hard to believe as a proper manifold because some fundamental theoretical issues arise regarding asymptotic flatness, singular behaviors, validity of energy conditions, etc., along with the stability issues. In [32, 33, 34, 35], some regular metric solutions were proposed for the multi-black hole configurations. In these models, electromagnetic fields or charged matter help to maintain the equilibrium between the black holes. However, these systems are generally unstable to small perturbations, which could cause them to merge or drift apart. Astorino and Vigano [29] were the first to propose an exact analytical solution for a configuration of two static black holes, where an external back-reacting gravitational field maintains the equilibrium between two sources without requiring additional matter fields. Most importantly, this metric is regular everywhere outside the event horizons, i.e., free from any conical and curvature singularities, providing a physical BBH metric solution. It is characterized by the masses and positions of the two black holes, external field parameters, and a gauge parameter. In this work, we investigate the transonic accretion flow around the primary black hole in the background of this metric. We find both the shock-free and shock-induced accretion solutions, and examine their corresponding disc properties, such as density distribution, temperature distribution, and emissivity, etc. Since the horizon of individual Schwarzschild black holes are distorted by their mutual interactions, the presence of the secondary black hole may influence the accretion properties of the CPD. To address this possibility, we explore the impact of binary parameters, specifically the mass ratio ($q$) of the two black holes and their separation ($z_{2}$) on the accretion solutions. We also examine the various shock properties and luminosity distributions of the CPD as functions of $q$ and $z_{2}$. These investigations show that the accretion properties of the CPD are nearly independent of the binary parameters, featuring earlier predictions based on different aspects of binary accretion [19].
@@ II Binary black hole spacetime
Astrophysical black holes are always embedded in external gravitational fields, which can be described by a series of multipole moments. It has been shown in [36] that any non-spherical matter distributions, such as accretion discs, galaxies, nebulae, etc., will generally produce a multipolar gravitational field. These external gravitational fields provide a physical mechanism to maintain the equilibrium between the black holes. The line element of a system of two Schwarzschild black holes immersed in a dipole-quadrupole external gravitational field can be expressed in cylindrical Weyl coordinates ($t,x,\phi,z$) as [29],
$\begin{split}ds^{2}&=-V(x,z)dt^{2}+x^{2}V^{-1}(x,z)d\phi^{2}\\ &+f(x,z)(dx^{2}+dz^{2}),\end{split}$ (1)
@@
In the above expressions, $W_{ij}=x^{2}+\mu_{i}\mu_{j}$, $Y_{ij}=(\mu_{i}-\mu_{j})^{2}$, and $\mu_{i}=\sqrt{x^{2}+(z-w_{i})^{2}}-(z-w_{i})$, where ($i,j$) take values from $1$ to $4$. The values of $w_{i}$ are chosen as $w_{1}=z_{1}-m_{1}$, $w_{2}=z_{1}+m_{1}$, $w_{3}=z_{2}-m_{2}$, and $w_{4}=z_{2}+m_{2}$. Here, the quantities $m_{i}$ and $z_{i}$ denote the mass and position of the $i$-th black hole, respectively. Additionally, $b_{1}$ and $b_{2}$ are the dipole and quadrupole momenta of the external gravitational field, respectively, while $C_{f}$ refers the gauge parameter.
@@
The static and axisymmetric metric (1) is an exact analytical solution of the vacuum Einstein’s equations in general relativity. It is important to note that the balance between two sources would be lost when the external gravitational field is absence ($b_{1}=b_{2}=0$). In such a scenario, Eq. (1) boils down to the usual Bach-Weyl metric [37, 38, 39], which also represents the double Schwarzschild black hole solution. In the binary system (1), two event horizons extend into the regions $w_{1}<z<w_{2}$ and $w_{3}<z<w_{4}$ when $x=0$, covering the curvature singularities of the spacetime. Moreover, it is not affected by any conical singularities along the $z$-axis under the regularity condition $f_{0}V\rightarrow 1$ as $x\rightarrow 0$ [40], which is valid everywhere outside the event horizons of individual sources. To satisfy the equation $f_{0}V=1$, we consider a particular model of the external gravitational background where the parameters $C_{f}$, $b_{1}$, and $b_{2}$ are expressed in terms of intrinsic black hole parameters ($m_{1}$, $m_{2}$, $z_{1}$, $z_{2}$) as [29],
@@
To analyze the event horizons of the two sources, we need to consider the near-horizon limit of metric (1). The near-horizon geometry of the primary black hole has been examined in [29]. In that study, the coordinates transformation $x=\sqrt{r(r-2m_{1})}\sin\theta$ (event horizon at $x_{\rm H}=0$) and $z=z_{1}+(r-m_{1})\cos\theta$ brings the metric (1) into a deformed Schwarzschild metric as $r$ approaches $2m_{1}$ (see Eq. (11) in [29]). So, it is evident that both the external fields and the secondary black hole can distort the primary black hole horizon. A similar description applies to the secondary black hole as well. Indeed, the binary black hole system (1) contains two distorted Schwarzschild black holes. The horizon area of primary black hole is calculated as [29],
@@
where $q$ ($=m_{2}/m_{1}$) is defined as the binary mass ratio. In this work, we use a unit system: $G=c=m_{1}=1$, where $G$ is the gravitational constant and $c$ is the speed of light. Also, we place the primary black hole at $z_{1}=0$. In Fig. 2, we illustrate the dependency of $A_{1}$ on mass ratio ($q$) and distance between two sources ($z_{2}$), where the variation of $A_{1}$ as a function of $q$ for different $z_{2}$ values has been displayed. The solid (black), dashed (blue), dotted (red), and dash-dotted (green) curves represent the obtained results for $z_{2}=300$, $400$, $500$, and $600$, respectively. We observed that for a given $z_{2}$, $A_{1}$ increases with $q$. However, for a fixed $q$, $A_{1}$ decreases with $z_{2}$. Therefore, $q$ and $z_{2}$ play opposite roles in determining the characteristics of $A_{1}$.

### v2

@@ Abstract
We study the general relativistic transonic accretion flow around the primary black hole, which forms the circumprimary disc (CPD), within a binary black hole (BBH) system. The BBH spacetime is characterized by the mass ratio ($q$) and the separation distance ($z_{2}$) between the two black holes. We numerically solve the radial momentum and energy equations to obtain the accretion solutions. It is observed that the CPD can exhibit shock solutions, which exist for a wide range parameter space spanned by flow specific angular momentum ($\lambda$) and energy ($E$). We find that the shock parameter space is modified by $q$ and $z_{2}$. Investigations show that $q$ and $z_{2}$ also affect various shock properties, such as density compression and temperature compression across the shock fronts. Moreover, we calculate the spectral energy distributions (SEDs) of the CPD and examine how the SEDs are modified by $q$ and $z_{2}$ for both shock-free and shock-induced accretion solutions. SED is found to be nearly independent of the binary parameters. We essentially show that although $q$ and $z_{2}$ alter the effective horizon area of the primary black hole located at the center of the CPD, they have a minimal impact on the dynamical and spectral properties of the accretion flow around the primary black hole.
@@ I Introduction
Binary black hole (BBH) systems are powerful multi-messenger astrophysical sources of gravitational waves (GWs) and electromagnetic (EM) waves. The formation scenarios of BBHs are not clearly understood yet. In this regard, various astrophysical channels have been proposed in the literature. Notable studies by Begelman et al. (1980) and Roos (1981) indicate that supermassive black hole binaries (SMBHBs) form as a result of galaxy mergers. On the other hand, stellar-mass BBHs are generated through dynamic encounters in the environments of dense star clusters Banerjee (2017); Rodriguez et al. (2018). Recently, it has been demonstrated that the gaseous dynamical friction can form a BBH system in the active galactic nucleus (AGN) discs DeLaurentiis et al. (2023).
@@
In addition to the understanding of the formation channels of BBHs, detecting them through observations is also a vital aspect in astrophysics. Meanwhile, LIGO and Virgo collaborations reveal that the gravitational interactions within BBHs can produce GWs Abbott et al. (2016), as predicted by general relativity, thereby providing direct evidence of stellar-mass BBHs through GWs. On the other hand, SMBHBs at sub-parsec separations are the strong GW sources Burke-Spolaor et al. (2019), which may be detectable by the Pulsar Timing Arrays Arzoumanian et al. (2020) or the future mission Laser Interferometer Space Antenna Amaro-Seoane et al. (2017); Mangiagli et al. (2022). Like AGNs, BBHs can radiate across the electromagnetic spectrum due to the accretion of matter from the surroundings. Therefore, in addition to GWs, analyzing EM signature from BBHs is also crucial for their study. In recent years, several EM observational works, including both the thermal and non-thermal emissions, have come out for the BBHs Roedig et al. (2014); D’Orazio et al. (2015); Graham et al. (2015a, b); Liu et al. (2019); Saade et al. (2020); Bogdanovic et al. (2022); Gutiérrez et al. (2022); Lee et al. (2024); Mondal et al. (2024). Indeed, these investigations provide the valuable information about the binary evolution, mass accretion rate, disc inclination angle, mass and the spin of the individual black holes, etc.
@@
The accretion flow onto a BBH system is qualitatively different from a single black hole accretion. For BBH accretion, a truncated disc surrounds the entire binary system Gutiérrez et al. (2024a), called the circumbinary disc (CBD). The gravitational potential of the binary determines the orbital structure of the CBD, while viscous stress enables accretion by transporting angular momentum outward Miranda et al. (2017). However, in the presence of a binary, the disc is truncated at a certain radius, which defines its inner boundary, where the outward tidal torques exerted by the binary balance the inward viscous torques Papaloizou and Pringle (1977); Lin and Papaloizou (1986); Artymowicz and Lubow (1994); Miranda et al. (2017); Heath and Nixon (2020). This balance typically occurs at Lindblad resonances Artymowicz and Lubow (1994), where disc material resonates with the binary’s orbital motion, resulting in the formation of a low-density cavity around the binary. The amount of matter that falls into the cavity depends on the disc scale height, thermodynamical properties (e.g., temperature and pressure gradients), and flow properties (e.g., viscosity) Artymowicz and Lubow (1996); D’Orazio et al. (2013); Heath and Nixon (2020); Tiede et al. (2020, 2022, 2025). The matter can flow into the cavity through narrow streams and begin to accrete onto the individual black holes, leading to the formation of a pair of mini-discs. Various recent studies have shown that this matter inflow can be highly variable and may exhibit characteristic periodicity depending on binary mass ratio and binary orbital parameters (e.g., eccentricity, orbital period) D’Orazio et al. (2013); Tiede et al. (2025). The mini discs surrounding the primary (more massive) and secondary (less massive) black holes are referred to as circumprimary discs (CPDs) and circumsecondary discs (CSDs), respectively. The presence of multiple discs (i.e., CBD, CPD, and CSD) complicates the study of BBH accretion compared to the accretion of individual black holes. Numerous theoretical models on BBHs accretion flow have been proposed in Newtonian gravity and general relativity (GR) (Sesana et al., 2011; Bowen et al., 2019; Lopez Armengol et al., 2021; Combi et al., 2022; Gutiérrez et al., 2022; Lee et al., 2024; Gutiérrez et al., 2024b, a, and references therein). All these models based on different physical conditions and explained the dynamics of CBD, mini discs, role of magnetic field, outflows or jets and EM spectral signature, etc., which indeed help to distinguish them from the ordinary AGNs. However, to the best of our knowledge, no body has conveyed the transonic acccretion flow for the BBH systems, where flow must satisfied the inner boundary condition set by the horizon Liang and Thompson (1980); Abramowicz and Zurek (1981). Toward this, for the first time we study the GR hydrodynamics of transonic accretion flow in the background of a BBH metric Astorino and Vigano (2021), which represents a system of two Schwarzschild black holes in an external gravitational field. We expect this study may provide the better understanding of the BBHs.
@@
Finding an exact analytical solution for binary or multiple black hole sources in GR is very challenging. Several attempts have been made to address this problem Kramer and Neugebauer (1980); Dietz and Hoenselaers (1985), where it often requires the extra matter fields, commonly interpreted as struts or cosmological strings, to support the gravitational attraction between the black holes. However, all of them are hard to believe as a proper manifold because some fundamental theoretical issues arise regarding asymptotic flatness, singular behaviors, validity of energy conditions, etc., along with the stability issues. In Majumdar (1947); Papapetrou (1945); Hartle and Hawking (1972); Emparan (2000), some regular metric solutions were proposed for the multi-black hole configurations. In these models, electromagnetic fields or charged matter help to maintain the equilibrium between the black holes. However, these systems are generally unstable to small perturbations, which could cause them to merge or drift apart. Astorino and Vigano (2021) were the first to propose an exact analytical solution for a configuration of two static black holes, where an external back-reacting gravitational field maintains the equilibrium between two sources without requiring additional matter fields. Most importantly, this metric is regular everywhere outside the event horizons, i.e., free from any conical and curvature singularities, providing a physical BBH metric solution. In their construction, they employed the Ernst framework, a formalism for generating exact stationary and axisymmetric solutions by reducing the Einstein field equations to a complex scalar equation for the Ernst potential. Within this framework, they used a technique in which the external gravitational field is modeled as a series of multipole moments. By appropriately choosing the multipole structure, they were able to achieve mutual equilibrium between the black holes against their gravitational attraction and also avoid any curvature singularities outside the horizons. In this way, they constructed a physically acceptable exact solution representing a realistic binary black hole configuration. It is characterized by the masses and positions of the two black holes, external field parameters, and a gauge parameter. In this work, we investigate the transonic accretion flow around the primary black hole in the background of this metric. We find both the shock-free and shock-induced accretion solutions, and examine their corresponding disc properties, such as density distribution, temperature distribution, and emissivity, etc. Since the horizon of individual Schwarzschild black holes are distorted by their mutual interactions, the presence of the secondary black hole may influence the accretion properties of the CPD. To address this possibility, we explore the impact of binary parameters, specifically the mass ratio ($q$) of the two black holes and their separation ($z_{2}$) on the accretion solutions. We also examine the various shock properties and luminosity distributions of the CPD as functions of $q$ and $z_{2}$. These investigations show that the accretion properties of the CPD are nearly independent of the binary parameters, featuring earlier predictions based on different aspects of binary accretion Lee et al. (2024).
@@ II Binary black hole spacetime
Astrophysical black holes are always embedded in external gravitational fields, which can be described by a series of multipole moments. It has been shown in de Castro and Letelier (2011) that any non-spherical matter distributions, such as accretion discs, galaxies, nebulae, etc., will generally produce a multipolar gravitational field. These external gravitational fields provide a physical mechanism to maintain the equilibrium between the black holes. The line element of a system of two Schwarzschild black holes immersed in a dipole-quadrupole external gravitational field can be expressed in cylindrical Weyl coordinates ($t,x,\phi,z$) as Astorino and Vigano (2021),
$\begin{split}ds^{2}&=-V(x,z)dt^{2}+x^{2}V^{-1}(x,z)d\phi^{2}\\ &+f_{0}(x,z)(dx^{2}+dz^{2}),\end{split}$ (1)
@@
In the above expressions, $W_{ij}=x^{2}+\mu_{i}\mu_{j}$, $Y_{ij}=(\mu_{i}-\mu_{j})^{2}$, and $\mu_{i}=\sqrt{x^{2}+(z-w_{i})^{2}}-(z-w_{i})$, where ($i,j$) take values from $1$ to $4$. The values of $w_{i}$ are chosen as $w_{1}=z_{1}-m_{1}$, $w_{2}=z_{1}+m_{1}$, $w_{3}=z_{2}-m_{2}$, and $w_{4}=z_{2}+m_{2}$. Here, the quantities $m_{i}$ and $z_{i}$ denote the mass and position of the $i$-th black hole, respectively. Additionally, $b_{1}$ and $b_{2}$ are the dipole and quadrupole moments in multipole expansion of the external gravitational field, respectively, while $C_{f}$ refers the gauge parameter.
@@
The static and axisymmetric metric (1) is an exact analytical solution of the vacuum Einstein’s equations in general relativity. It is important to note that the balance between two sources would be lost when the external gravitational field is absence ($b_{1}=b_{2}=0$). In such a scenario, Eq. (1) boils down to the usual Bach-Weyl metric Bach and Wevl (1922); Israel and Khan (1964); Costa and Perry (2000), which also represents the double Schwarzschild black hole solution. In the binary system (1), two event horizons extend into the regions $w_{1}<z<w_{2}$ and $w_{3}<z<w_{4}$ when $x=0$, covering the curvature singularities of the spacetime. Moreover, it is not affected by any conical singularities along the $Z$-axis under the regularity condition $f_{0}V\rightarrow 1$ as $x\rightarrow 0$ Alekseev and Belinski (2019), which is valid everywhere outside the event horizons of individual sources. To satisfy the equation $f_{0}V=1$, we consider a particular model of the external gravitational background where the parameters $C_{f}$, $b_{1}$, and $b_{2}$ are expressed in terms of intrinsic black hole parameters ($m_{1}$, $m_{2}$, $z_{1}$, $z_{2}$) as Astorino and Vigano (2021),
@@
To analyze the event horizons of the two sources, we need to consider the near-horizon limit of metric (1). The near-horizon geometry of the primary black hole has been examined in Astorino and Vigano (2021). In that study, the coordinates transformation $x=\sqrt{r(r-2m_{1})}\sin\theta$ (event horizon at $x_{\rm H}=0$) and $z=z_{1}+(r-m_{1})\cos\theta$ brings the metric (1) into a deformed Schwarzschild metric as $r$ approaches $2m_{1}$. The expression of this deformed Schwarzschild metric is given by Astorino and Vigano (2021),
$\begin{split}ds^{2}&\simeq h_{0}(\theta)\left[-\left(1-\frac{2m_{1}}{r}\right)e^{F_{1}(\theta)}dt^{2}+\frac{D^{2}e^{F_{2}(\theta)}}{1-\frac{2m_{1}}{r}}dr^{2}\right]\\ &+(2m_{1})^{2}\left[h_{0}(\theta)D^{2}e^{F_{2}(\theta)}d\theta^{2}+\frac{\sin^{2}\theta}{h_{0}(\theta)e^{F_{1}(\theta)}}d\phi^{2}\right],\end{split}$ (4)
where
$\begin{split}h_{0}(\theta)&=\frac{m_{1}\cos\theta+m_{2}+z_{1}-z_{2}}{m_{1}\cos\theta-m_{2}+z_{1}-z_{2}},\\ D&=\frac{m_{1}+m_{2}-z_{1}+z_{2}}{m_{1}-m_{2}-z_{1}+z_{2}},\\ F_{1}(\theta)&=2\left[b_{1}+b_{2}(z_{1}+m_{1}\cos\theta)\right](z_{1}+m_{1}\cos\theta),\\ F_{2}(\theta)&=2b_{1}[m_{1}\cos\theta-2m_{1}-4m_{2}-z_{1}]\\ &+2b_{2}[m_{1}^{2}\cos^{2}\theta+2m_{1}z_{1}\cos\theta\\ &-2m_{1}^{2}-z_{1}^{2}-4m_{1}z_{1}-8m_{2}z_{2}].\end{split}$ (5)
Note that when both the external field and the second black hole vanish, one finds from Eq. (5) that $h_{0}(\theta)=1$, $D=1$, $F_{1}(\theta)=0$, and $F_{2}(\theta)=0$. Therefore, Eq. (4) reduces to the usual Schwarzschild black hole of mass $m_{1}$. Also, from Eq. (4) it is evident that both the external fields and the secondary black hole can distort the primary black hole horizon. A similar description applies to the secondary black hole as well. Indeed, the binary black hole system (1) contains two distorted Schwarzschild black holes. The horizon area of primary black hole is calculated as Astorino and Vigano (2021),
@@
where $q$ ($m_{2}/m_{1}$) is defined as the binary mass ratio. In this work, we use a unit system: $G=c=m_{1}=1$, where $G$ is the gravitational constant and $c$ is the speed of light. Also, we place the primary black hole at $z_{1}=0$. In Fig. 2, we illustrate the dependency of $A_{1}$ on mass ratio ($q$) and distance between two sources ($z_{2}$), where the variation of $A_{1}$ as a function of $q$ for different $z_{2}$ values has been displayed. The solid (black), dashed (blue), dotted (red), and dash-dotted (green) curves represent the obtained results for $z_{2}=300$, $400$, $500$, and $600$, respectively. We observed that for a given $z_{2}$, $A_{1}$ increases with $q$. However, for a fixed $q$, $A_{1}$ decreases with $z_{2}$. Therefore, $q$ and $z_{2}$ play opposite roles in determining the characteristics of $A_{1}$.
In the introduction section, we provide a detailed discussion on how the authors in Astorino and Vigano (2021) proposed a static BBH metric by considering phenomenologically realistic scenarios as much as possible. In that study, to maintain the equilibrium between the black holes against the gravitational attraction between them, the authors did not include singular matter (cosmological strings) or charged matter as these matter fields routinely introduce the singularities outside the event horizons. Instead, they considered the black holes to be immersed in an external gravitational field. Specifically, they included only the dipole and quadrupole moments in the multipolar expansion of the external field, as these are sufficient to stabilize the BBH system. Also, the multipolar expansion is a key approach for regularizing the nodal singularity, due to it’s foundation in Ernst’s insight. However, as the masses of the black holes continuously increase due to the accretion of matter from the surroundings, it becomes challenging for the external fields to maintain stability in the system. When the black holes masses grow sufficiently, such that the gravitational attraction between them exceeds the repulsion from the external fields, the BBH system will collapse into a single black hole. However, the merging of black holes is an adiabatic process, i.e., it will take a very long time ($\sim 10^{7}-10^{9}$ yr) to occur, as the mass accretion rates of supermassive black holes are very small Yarza et al. (2020). Over this extended timescale, the system is expected to remain nearly stable, with the two black holes gradually drifting toward each other in a manner that is nearly negligible. Also, in our work, we focus on scenarios with large binary separations. Therefore, when studying the accretion physics over such a prolonged timescale, the metric we have used is expected to closely resemble the nearly static configuration of a BBH system. However, in a real BBH accretion scenario, applying the metric (1) would be challenging, as the system is inherently dynamic. Since the primary goal of this work is to investigate how the presence of a secondary black hole influences the event horizon of the primary black hole and, consequently, its accretion properties, we adopt this simplified BBH metric.